# The Tautochrone of Huygens and Abel: From Constructive Geometry to Fractional Calculus

### v1

@@ 3.1 Abel’s Solution: Remarkable Theorem
Abel then considers a problem more general than the classical tautochrone, which we now briefly summarise in order to follow the core ideas and consequences of his approach. Consider Fig. 2, which illustrates the motion of a particle along an arc, such as the one described by the curve CA. Let CB be a horizontal line and AB perpendicular to CB. The curve AM is described in Cartesian coordinates, where $AP=x$ and $PM=y$. We also define $AB=a$ and denote by $AM=s$ the arc length along the curve, which is initially unknown.
The particle begins its motion from rest at point C and moves along the curve CA. The time $T$ required to traverse the arc depends both on the shape of the curve and on the distance $a$. The problem is thus to determine the form of the curve KCA such that the time of descent $T$ is a given function of $a$, for instance, $T=\psi(a)$. We use today’s notation; the notation employed by Abel is such that a function $\psi(a)$ is denoted by $\psi\,a$.
@@
Figure 2: The illustrative picture appearing in the French translation of the Abel’s 1823 paper [17].
After some mechanical considerations, Abel found for this problem the following equation defining the time involved in the motion of the particle:
$\psi(a)=\int_{x=0}^{x=a}\frac{ds}{\sqrt{a-x}},$ (8)
in which $s=s(x)$ is the unknown function. However, instead to handle with it directly, Abel considers the more general equation:
$\psi(a)=\int_{x=0}^{x=a}\frac{ds}{(a-x)^{n}},$ (9)
where $n$ has to be less than 1 to prevent the infinite integral between two limits; $\psi(a)$ is an arbitrary function that is not infinite, when $a=0$.
To proceed, let us recall that the problem is to find the curve $s(x)$ which Abel searches in the form of a power series
$s(x)=\sum_{m}a^{(m)}x^{m}.$ (10)
After term-by-term operations with the series, and after using the properties of the gamma function — by quoting in this regard the work of A. M. Legendre —, and after some pages of manipulations with the power series, he arrives at the following solution to Eq. (9), called by him as a “remarkable theorem” [20]:
$s(x)=\frac{\sin n\pi}{\pi}x^{n}\int_{0}^{1}\frac{\psi(xt)dt}{(1-t)^{1-n}}.$ (11)
In a few words, the remarkable result is that given an integral like the one in Eq. (9), the curve $s(x)$ is given by Eq. (11). After examining the obtained formulas from other perspectives, Abel mentions that $s(x)$ can be expressed in a different way, namely:
$s(x)=\frac{1}{\Gamma(1-n)}\int^{n}\psi(x)dx^{n}=\frac{1}{\Gamma(1-n)}\frac{d^{-n}}{dx^{-n}}\psi(x).$ (12)
As underlined in Ref. [20], it is easy to recognize, behind the archaic notation he used, that Abel presents two expressions for the fractional-order integral: a derivative of negative order and a symbol that later appears in Liouville’s works for the integral of order $n$: $\int^{n}\psi(x)dx^{n}$ (See Refs. [21, 22, 23]).
Subsequently, Abel considers the case $n=1/2$, in order to obtain from the integral
$\psi(a)=\int_{x=0}^{x=a}\frac{ds}{\sqrt{a-x}},$ (13)
the desired curve $s(x)$, when the time is given by $psi(a)$, i.e.,
$s(x)=\frac{1}{\sqrt{\pi}}\frac{d^{-1/2}}{dx^{-1/2}}\psi(x)=\frac{1}{\sqrt{\pi}}\int^{1/2}\psi(x)dx^{1/2},$ (14)
from which one obtains:
$\psi(x)=\sqrt{\pi}\frac{d^{1/2}}{dx^{1/2}}s(x).$ (15)
For our analysis, this is a crucial point in the paper because Abel inverts the fractional integral by using a fractional differentiation on both sides of the equation, thus assuming (and understanding) that fractional order integration and differentiation are mutually inverse procedures. He affirms indeed [17]:
@@
If the equation of a curve is $s=\psi(x)$, the time a body takes to travel along an arc of the curve whose height is $a$ is $\sqrt{\pi}\frac{d^{1/2}\psi(a)}{da^{1/2}}$.
To conclude this incursion for the original approach of Abel, we remark for future purposes that Eq. (9) could be rewritten as:
$\psi(t)=\int_{0}^{t}\frac{s^{\prime}(x)}{(t-x)^{n}}dx,$ (16)
which, apart from a constant coefficient, may be identified with the well-known Caputo’s time-fractional derivative of real order to be introduced later [24, 25]. In this perspective, the solution proposed by Abel is the inverse operation of the fractional integral of the same arbitrary order $n$ [20].
3.2 Abel’s Isochrone: Explicit Solution
To appreciate the result obtained by Abel in more simple terms, let us first reformulate the physical problem of a particle of mass $m$ sliding the curve shown in Fig. 3. It is released from the rest at the position $y=y_{0}$ and slides under the influence of the gravity, along a frictionless wire.
@@ 3.2 Abel’s Isochrone: Explicit Solution
The problem is to determine the curve $s(y)$, followed by the particle, such that the time employed to slide down to the lowest point under gravity is independent of its initial position ($x_{0},y_{0}$) on the curve. If the mass is released from $y=y_{0}>0$, then its speed at $y$ will be given by
$\frac{1}{2}m\left(\frac{ds}{dt}\right)^{2}=mgy_{0}-mgy=mg(y_{0}-y),$ (17)
from which we obtain:
$dt=\pm\frac{ds}{\sqrt{2g(y_{0}-y)}}\quad\longrightarrow\quad dt=-\frac{1}{\sqrt{2g(y_{0}-y)}}\,\frac{ds}{dy}\,dy,$ (18)
because $ds/dy>0$, but $dy/dt<0$. The time of descent is then
$T(y_{0})=\int_{0}^{y_{0}}\frac{1}{\sqrt{2g(y_{0}-y)}}\,\frac{ds}{dy}\,dy.$ (19)
We observe that (19) was obtained by invoking only the conservation of the mechanical energy, Eq. (17). It is an integral equation to be solved in order to get $f(y)=ds/dy$. Let us change a little the notation ($y_{0}\to y,\,y\to z$) to search for a curve $s(z)$ such that
$\sqrt{2g}T(y)=\int_{0}^{y}\frac{1}{\sqrt{y-z}}\frac{ds}{dz}\,dz=k(y)=\textrm{constant}.$ (20)
The formal problem is then to find $f(z)=ds/dz$, given $k(y)$:
$k(y)=\int_{0}^{y}\frac{f(z)}{\sqrt{y-z}}dz=\int_{0}^{y}f(z)K(y,z)dz,$ (21)
which may be recognized as a convolution between $f(z)$ and a kernel $K(y,z)=(y-z)^{-1/2}$. Equation (21) is an integral equation for $f(z)$ and may be solved by taking the Laplace transform of its both sides. Remembering that the Laplace transform is defined as
${\mathcal{L}}\{f(t);s\}:=F(s)=\int_{0}^{\infty}e^{-st}f(t)dt,\quad\quad s\in\mathbb{C},$ (22)
where $\mathbb{C}$ stands for the set of complex numbers, after its computation, we obtain
$\frac{k(y)}{s}=F(s)\frac{\sqrt{\pi}}{s^{1/2}}\quad\Longleftrightarrow\quad F(s)=\frac{c_{1}}{s^{1/2}},$ (23)
where $c_{1}=k(y)/\sqrt{\pi}$ is a constant. The inverse transform of $F(s)$ is
$f(z)=cz^{-1/2},$ (24)
where $c$ is another constant. Remember that $f(z)=ds/dz$ or $f(y)=ds/dy$. Thus,
$f(y)=\frac{ds}{dy}=\left[1+\left(\frac{dx}{dy}\right)^{2}\right]^{1/2}=cy^{-1/2}.$ (25)
By putting $y=c^{2}\sin^{2}(1/2\theta)$, a short calculation yields
$x=\frac{1}{2}c^{2}(\theta+\sin\theta)\quad{\textrm{and}}\quad y=\frac{1}{2}c^{2}(1-\cos\theta),$ (26)
which are the parametric equations of a cycloid [26].
3.3 The Birth of Fractional Calculus

### v2

@@ III.1 Abel’s Solution: A Remarkable Theorem
He considers a generalization of the tautochrone problem, illustrated in Fig. 2. A particle descends from rest at point C along an unknown curve CA. Let $AB=a$ be the vertical height, and denote by $s(x)$ the arc length as a function of the horizontal coordinate $x=AP$. The descent time $T$ is prescribed as a given function $T=\psi(a)$. The problem is to determine the curve $s(x)$ that yields this time.
@@
Figure 2: The illustrative picture appearing in the French translation of the Abel’s 1823 paper Abel1823 .
Abel derives the integral equation
$\psi(a)=\int_{x=0}^{a}\frac{ds}{\sqrt{a-x}},$ (4)
where $ds=s^{\prime}(x)\mskip 3.0mudx$. Instead of solving this specific case directly, he treats the more general equation
$\psi(a)=\int_{0}^{a}\frac{ds}{(a-x)^{n}},\qquad 0<n<1,$ (5)
with $\psi(a)$ finite at $a=0$.
After a series of manipulations involving power–series expansions and properties of the gamma function, Abel arrives at what he calls a “remarkable theorem” Podlubny-Magin :
$s(x)=\frac{\sin n\pi}{\pi}\mskip 3.0mux^{n}\int_{0}^{1}\frac{\psi(xt)\mskip 3.0mudt}{(1-t)^{1-n}}.$ (6)
In modern notation, this solution can be written as
$s(x)=\frac{1}{\Gamma(1-n)}\frac{d^{-n}}{dx^{-n}}\mskip 3.0mu\psi(x),$ (7)
which Abel recognizes as a fractional-order integral — an operation inverse to fractional differentiation. Indeed, he later writes:
@@
If the equation of a curve is $s=\psi(x)$, the time a body takes to travel along an arc of the curve whose height is $a$ is $\sqrt{\pi}\mskip 3.0mu\dfrac{d^{1/2}\psi(a)}{da^{1/2}}$.
For the tautochrone ($n=1/2$), the solution reduces to
$\psi(x)=\sqrt{\pi}\mskip 3.0mu\frac{d^{1/2}}{dx^{1/2}}\mskip 3.0mus(x),$ (8)
which inverts the original integral. Crucially, Abel treats fractional integration and differentiation as mutually inverse procedures — a conceptual leap that foreshadows the later development of fractional calculus (see Refs. Liouville1832b ; Liouville1832a ; Liouville1834 ).
Equation (5), rewritten as
$\psi(t)=\int_{0}^{t}\frac{s^{\prime}(x)}{(t-x)^{n}}\mskip 3.0mudx,$ (9)
is — up to a constant factor — precisely what would later be called the Caputo fractional derivative of order $n$ Caputo1967 ; Caputo1969 . Abel’s solution therefore represents the first systematic inversion of such an integral equation, establishing a direct link between the tautochrone problem and the birth of fractional calculus.
III.2 Abel’s Isochrone: Explicit Solution
To derive Abel’s result in simpler terms, consider a particle of mass $m$ sliding from rest at height $y_{0}$ along a frictionless curve $s(y)$, as shown in Fig. 3. The descent time must be independent of $y_{0}$.
@@ III.2 Abel’s Isochrone: Explicit Solution
Conservation of energy gives $v=\sqrt{2g(y_{0}-y)}$, so that
$dt=-\frac{ds}{\sqrt{2g(y_{0}-y)}}=-\frac{1}{\sqrt{2g(y_{0}-y)}}\frac{ds}{dy}\mskip 3.0mudy.$
The total descent time is therefore
$T(y_{0})=\int_{0}^{y_{0}}\frac{1}{\sqrt{2g(y_{0}-y)}}\mskip 3.0mu\frac{ds}{dy}\mskip 3.0mudy.$ (10)
Equation (10) is an integral equation for $f(y)=ds/dy$. Writing $y_{0}\to y$ and $y\to z$, and noting that tautochrony requires $T(y)=\text{constant}/\sqrt{2g}$, we obtain
$k=\int_{0}^{y}\frac{f(z)}{\sqrt{y-z}}\mskip 3.0mudz,$ (11)
where $k$ is constant. This is a convolution with kernel $K(y,z)=(y-z)^{-1/2}$.
The equation is readily solved via Laplace transforms. Using
$\mathcal{L}\left\{\frac{1}{\sqrt{y}};s\right\}=\sqrt{\frac{\pi}{s}},$
the transform of (11) yields
$\frac{k}{s}=F(s)\sqrt{\frac{\pi}{s}}\quad\Longrightarrow\quad F(s)=\frac{c}{s^{1/2}},$
with $c=k/\sqrt{\pi}$. Inverting gives
$f(y)=\frac{ds}{dy}=c\mskip 3.0muy^{-1/2}.$
Since $ds/dy=\sqrt{1+(dx/dy)^{2}}$, we have
$1+\left(\frac{dx}{dy}\right)^{2}=\frac{c^{2}}{y}.$
Setting $y=\frac{c^{2}}{2}(1-\cos\theta)$ leads to the parametric equations
$x$ $=\frac{c^{2}}{2}(\theta+\sin\theta),$ (12)
$y$ $=\frac{c^{2}}{2}(1-\cos\theta),$ (13)
which define a cycloid Sokolnikoff1958 . Thus Abel’s inversion of the integral equation recovers the same curve that Huygens had verified geometrically a century and a half earlier.
III.3 The Birth of Fractional Calculus